OpenStudy (gabylovesyou):

Which of the following is not an equivalent form of the compound inequality x + 18 /= 5 A number line with a closed circle on -13, shading to the left, and a closed circle on -8, shading to the right. 10 >/= x + 18 >/= 5 -13

OpenStudy (anonymous):

a NOT question again.... you have wierd problems gaby....

OpenStudy (gabylovesyou):

lmaooooo

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

since this is an "AND" statement, the solution set is an intersection of two sets...

OpenStudy (phi):

I would simplify the first statement x + 18 <= 10 and x + 18 >= 5 and see if it matches the last statement -13 <= x <= -8 I would "unglue" the 2nd statement 10 >= x + 18 >= 5 and see if it matches the question then I would plot (sketch) the graphs, and ask if they match the problem. one of the four should not match. That is your answer